1.108.2 Find Linux documentation on the Internet Weight 3

Linux Professional Institute Certification — 102

Geoffrey Robertson ge@ffrey.com

               Nick Urbanik nicku@nicku.org

2005 July

Description of Objective

Candidates should be able to find and use Linux documentation. This objective includes using Linux documentation at sources such as the Linux Documentation Project (LDP), vendor and third-party websites, newsgroups, newsgroup archives, and mailing lists.

Key files, terms, and utilities include:

not applicable

Documentation 2.108.1 Use and Manage Local System Documentation 2.108.2

Find Linux Documentation on the Internet[] 2.108.5 Notify Users on System Related Issues

\underline{2}

The End License Of This Document